Individual retirement accounts (IRAs) were established by the U.S. government to encourage saving. An individual who deposits part of current earnings in an IRA does not have to pay income taxes on the earnings deposited, nor are any income taxes charged on the interest earned by the funds in the IRA. However, when the funds are withdrawn from the IRA, the full amount withdrawn is treated as income and is taxed at the individual’s current income tax rate. In contrast, an individual depositing in a non-IRA account has to pay income taxes on the funds deposited and on interest earned in each year but does not have to pay taxes on withdrawals from the account. Another feature of IRAs that is different from a standard savings account is that funds deposited in an IRA cannot be withdrawn prior to retirement, except upon payment of a substantial penalty.

a. Greg, who is five years from retirement, receives a $10,000 bonus at work. He is trying to decide whether to save this extra income in an IRA account or in a regular savings account. Both accounts earn 6 percent nominal interest, and Greg is in the 30 percent tax bracket in every year (including his retirement year). Compare the amounts that Greg will have in five years under each of the two saving strategies, net of all taxes. Is the IRA a good deal for Greg?

Instructions: Round your responses to the nearest dollar.

If Greg invests in the IRA, his net value (after taxes) five years from now will be: $.

If Greg invests in the normal savings account, his net value (after taxes) five years from now will be: $.

Greg will be better off if he invests in the (Click to select)regular savings accountIRA.


b. Would you expect the availability of IRAs to increase the amount that households save in light of:

(1) the response of saving to changes in the real interest rate? (Click to select)YesNo.

(2) psychological theories of saving? (Click to select)YesNo.
Answer #1

Investment in IRA. By using the formula of compound interest, we can calculate the amount ater 5 years:

A = P (1+r)n = 10000 (1+.06)5 = 13382.25.

Income tax on total amount @ 30% = 4014.67. Therefore future value net of taxes = 9367.58

Investment in Savings Account:

Op bal / Investment Closing Year Interest Tax investment 10,000.00 7,420.00 7,731.64 8,056.37 8,394.74 600.00 445.20 463.90 4

Thus it is clear from the above calculations, that the net amount after tax is more when invested in IRA. Thus Greg will be better off if he invests in IRA.